5th grade:
2x - y = 12
What does x = and y =

Answer by checkley75(3666)

2x-y=12
2x=y+12
x=y/2+12/2
x=y/2+6
----------------------
2x-y=12
-y=-2x+12
y=2x-12
y=2(x-6)
----------------------
there is an infinite number of solution for x & y.
if x=6 then 6=y/2+6
y/2=6-6
y/2=0
y=0
---------------------
if y=12 then 12=2(x-6)
12=2x-12
2x=24
x=24/2
x=12
---------------------
try some other x or y values to prove the point.